# Supplementary Materials

### v1

@@ Abstract
Learning Equivariant Functions via Quadratic Forms
Pavan Karjol1          Vivek V Kashyap1          Rohan Kashyap2          Prathosh A P1
1Department of Electrical Communication Engineering, Indian Institute of Science, Bengaluru, Karnataka 2Computer Science, Carnegie Mellon University, Pittsburgh, Pennsylvania, USA
@@ 1 Introduction
Symmetry serves as a fundamental principle in modern deep learning, providing powerful inductive biases that enhance model generalization and sample efficiency Cohen (2013); Cohen and Welling (2014); Finzi et al. (2021); Dehmamy et al. (2021). When the symmetry group underlying a system is known a priori, this symmetry can be explicitly encoded into the neural network architecture through invariant or equivariant operations. This architectural integration ensures the model maintains equivariance with respect to the predefined group transformations present in the data manifold. The classical example is the Convolutional Neural Networks (CNNs), which maintain translational equivariance through weight sharing across spatial dimensions.
Recent works have explored two primary approaches to incorporating symmetries: augmenting training data using symmetry structures (Benton et al., 2020) and designing architectures with intrinsic symmetry properties (Zaheer et al., 2017; Kicki et al., 2020). (Benton et al., 2020) propose a data augmentation approach by parameterizing the data distribution while (Zhou et al., 2020) introduce a meta-learning framework through reparameterization of the network layers. In our work, we present a framework for learning group-equivariant functions through quadratic forms, emphasizing applications to orthogonal groups. Our analysis proves that these functions admit a canonical decomposition into two fundamental components: scale-invariant and norm-invariant terms.
Although substantial progress has been achieved in harnessing symmetries across various domains—including images (Anselmi et al., 2019; Kondor and Trivedi, 2018), graphs (Han et al., 2022), and dynamical systems such as partial differential equations (PDEs) (Horie and Mitsume, 2023)—significant challenges remain. A central difficulty in integrating these group symmetries within model architectures is the often-required explicit, a priori knowledge of the underlying symmetry. Furthermore, existing models frequently suffer from high computational demands, stemming from requirements such as null-space computations and implementation of exponential maps. This complexity limits their scalability and flexibility. Additionally, in many practical applications, the group action on the data may not be immediately apparent, making it difficult to effectively leverage these symmetries in current group-equivariant models.
Consequently, existing models are constrained in their ability to effectively manage complex group symmetries, especially within real-world data contexts. This limitation is similarly observed in dynamical systems, where simple symmetries in a low-dimensional state space can appear as highly non-linear transformations when projected into high-dimensional measurements. In this work, we propose a novel framework for learning equivariant functions, which integrates two invariant functions and is applicable regardless of whether the underlying symmetry group is known or unknown, given certain constraints.
Our approach focuses on orthogonal groups and their inherent geometric properties. Specifically, we are motivated by a fundamental property of the standard orthogonal group $O(n)$: its preservation of the Euclidean norm, which leads to orbits defined by distinct norm values. In addition, we also observe scale invariance in the case $O(n)$-equivariant functions. We demonstrate that this characteristic extends to other orthogonal groups preserving an arbitrary quadratic form, $x^{\top}Ax$, and we exploit these geometric correspondences to construct a systematic framework for symmetry discovery via neural networks. Furthermore, we explore the Lorentz group, which preserves the Minkowski norm and embodies the essential symmetry underlying the spacetime structure in physics. In localized regions of spacetime, where gravitational effects are negligible, physical laws exhibit Lorentz invariance in a manner consistent with special relativity.
In Section 4, we provide a comprehensive mathematical framework for our proposed method, particularly with respect to orthogonal groups, establishing its scalability and efficiency as a model for real-world applications.
@@ 1.1 Contributions
  • We establish that any $G$-equivariant function $f:X\rightarrow\mathbb{R}^{m}$, with $X\subset\mathbb{R}^{n}$ and $G$ being the orthogonal group preserving the quadratic form $x^{T}Ax$, can be represented as the product of two invariant functions—one that is norm-invariant and another that is scale-invariant—with respect to the underlying quadratic form. Here, the product is to be understood in the sense of the group action.
  • We extend this representation to functions defined on tuples of input vectors via a diagonal (product) group action on the product space $X_{p}=X\times\cdots\times X$. In this extended setting, the equivariant function is decomposed into an angular component—extracted solely from the normalized first vector—and a scale-invariant component that depends on the full Gram matrix of the tuple.
  • Leveraging this canonical decomposition into scale- and norm-invariant functions, we introduce a framework for symmetry discovery in cases where the underlying symmetry group is unknown. In this framework, uncovering the symmetry in the data corresponds to learning the appropriate quadratic form, which enables us to adaptively identify the symmetries present within a given dataset.
  • We validate our proposed method across several tasks that leverage invariance and equivariance properties. In particular, we demonstrate its effectiveness on polynomial regression, top quark tagging classification (which exploits Lorentz invariance), and the prediction of moment of inertia matrices (an $O(n)$-equivariant task).
@@ 2.2 Automatic Symmetry Discovery
Advances have also been made in the area of automatic symmetry discovery for neural networks. (Dehmamy et al., 2021) introduced Lie algebra convolution networks (L-Conv), achieving equivariance to continuous groups by leveraging Lie algebras. Complementary to this, (Benton et al., 2020) developed parameterized distributions over training data augmentations, while (Zhou et al., 2020) proposed a meta-learning approach that reparameterizes network layers to address symmetry adaptivity. Focusing on invariant representations, (Anselmi et al., 2019) explored regularization techniques for unsupervised orbit learning. (Yang et al., 2023) introduced LieGAN, a generative adversarial framework for symmetry discovery that leverages the principle of distribution preservation under group transformations.
More recently, (Yang et al., 2024) proposed Latent LieGAN, which approximates non-linear symmetries via non-linear group actions, particularly suited for high-dimensional dynamical systems. Additionally, (Karjol et al., 2023) introduced a structural decomposition approach for identifying locally symmetric, dihedral, and cyclic subgroups through multi-armed bandit algorithms.
@@ 3 Preliminaries
A group $(G,\cdot)$ is a set G equipped with a binary operation $\cdot:G\times G\to G$ satisfying the following properties:
  • Closure: For all $a,b\in G$, $a\cdot b\in G$.
  • Associativity: For all $a,b,c\in G$, $(a\cdot b)\cdot c=a\cdot(b\cdot c)$.
  • Identity element: There exists an element $e\in G$ such that for every $a\in G$, $e\cdot a=a\cdot e=a$.
  • Inverse element: For each $a\in G$, there exists an element $b\in G$ (often denoted $a^{-1}$) such that $a\cdot b=b\cdot a=e$.
@@
A group action of a group $G$ on a set $X$ is a function $G\times X\to X,\quad(g,x)\mapsto g\cdot x,$ such that for all $g,h\in G$ and $x\in X$:
  • Identity: $e\cdot x=x$, where $e$ is the identity element of $G$.
  • Compatibility: $(g\cdot h)\cdot x=g\cdot(h\cdot x)$.
@@
Given a group action of $G$ on a set $X$ the orbit of an element $x\in X$ under $G$ is the set:
@@
The general linear group $\mathrm{GL}(n,\mathbb{R})$ is the group of all $n\times n$ invertible matrices with entries from real numbers with group operation given by matrix multiplication.
Definition 3.5 (Invariant Function).
An invariant function under a group action is a function $f:X\to Y$ that remains unchanged when the group acts on its input. Formally, if $G$ acts on $X$, then $f$ is invariant if $f(g\cdot x)=f(x)$ for all $g\in G$ and $x\in X$. Invariance often signifies that $f$ depends only on certain properties that are preserved by the group action.
@@
An equivariant function between two sets with group actions is a function that respects the group action. If a group $G$ acts on both sets $X$ and $Y$, then a function $f:X\to Y$ is equivariant if $f(g\cdot x)=g\cdot f(x)$ for all $g\in G$ and $x\in X$. Equivariance means $f$ “commutes” with the group action and transforms in a manner consistent with the structure of $G$. Invariant function is a special case of Equivariant Function where $g\cdot f(x)=f(x),\ \quad\forall g\in G$.
@@ 4 Proposed Method
4.1 Quadratic form
Quadratic form is a type of mathematical expression in which a function of several variables involves terms of degree two (or order two). Let $A\in\mathbb{R}^{n\times n}$ be an $n\times n$ matrix and $x\in\mathbb{R}^{n}$. In general, a quadratic form in $x\in\mathbb{R}^{n}$ can be written as,
$q\mathopen{}\left({x;A}\,\negthickspace\right)\mathclose{}:=x^{T}Ax$ (2)
Each quadratic form defined on $\mathbb{R}^{n}$ is uniquely determined by a symmetric matrix. This is because, we can write,
$q(x;A)=x^{T}Ax=\frac{1}{2}\mathopen{}\left[{x^{T}\mathopen{}\left({A+A^{T}}\,\negthickspace\right)\mathclose{}x}\,\negthickspace\right]\mathclose{},\quad\forall x\in X$ (3)
Here, the $\frac{1}{2}\mathopen{}\left({A+A^{T}}\,\negthickspace\right)\mathclose{}$ is the unique symmetric matrix for the quadratic form $q\mathopen{}\left({.;A}\,\negthickspace\right)\mathclose{}$. Hence, we can simply assume that $A$ is symmetric matrix in any quadratic form $q(x;A)$.
4.1.1 Diagonalizing the Quadratic Form
Given a quadratic form $q(x;A)=x^{T}Ax$, where $A$ is an $n\times n$ symmetric matrix, we aim to simplify this form by transforming it into a diagonal form.
Since $A$ is symmetric, it can be diagonalized by an orthogonal matrix $U$. This means there exists an orthogonal matrix $U$ such that:
@@ 4.1.1 Diagonalizing the Quadratic Form
where $D$ is a diagonal matrix containing the eigenvalues of $A$.
Now, we can express the quadratic form in terms of this diagonalization as follows:
$q\mathopen{}\left({x;A}\,\negthickspace\right)\mathclose{}=x^{T}\mathopen{}\left({U^{T}DU}\,\negthickspace\right)\mathclose{}x=\mathopen{}\left({Ux}\,\negthickspace\right)\mathclose{}^{T}D\mathopen{}\left({Ux}\,\negthickspace\right)\mathclose{}=y^{T}Dy,$ (5)
where $y=Ux$. Thus, learning the quadratic form $q(x;A)$ reduces to learning an orthogonal matrix $U$ and a diagonal matrix $D$.
4.2 Orthogonal Group: $O(n;q)$
Consider the set,
$G:=\mathopen{}\left\{{U\in GL\mathopen{}\left({n;\mathbb{R}}\,\negthickspace\right)\mathclose{}:U^{T}AU=A}\,\negthickspace\right\}\mathclose{},$ (6)
which can easily be shown to be a Lie subgroup of $GL(n;\mathbb{R})$. Such groups are called orthogonal groups, denoted as $O(n;q)$, and they preserve the quadratic form $q(x;A)=x^{T}Ax$, meaning that $q(x;A)=q(Rx;A)$ for all $R\in O(n;q)$.
Examples of orthogonal groups include the standard orthogonal group that preserves the Euclidean norm $x^{T}x$, and the Lorentz group, which preserves the Minkowski norm $x^{T}\eta x$ where $\eta=\operatorname{diag}\left(1,-1,-1,-1\right)$, among others.
@@ 4.3 Diagonal group action
Let $A$ be a non-singular $n\times n$ real matrix and define Let
$G:=\{g\in\mathrm{GL}(n,\mathbb{R}):g^{T}A\,g=A\}$
be the orthogonal group preserving the quadratic form $x^{T}Ax$. Let $X\subset\mathbb{R}^{n}$ be a $G$-invariant set. For any positive integer $p$, consider the diagonal action of $G$ on
$X_{p}:=\underbrace{X\times X\times\cdots\times X}_{p\text{ times}},$
defined by
Then every $G$-equivariant function
$f:X_{p}\to\mathbb{R}^{m}$
can be decomposed as
@@
$\operatorname{Gram}(x_{1},\ldots,x_{p})=\begin{pmatrix}x_{1}^{T}Ax_{1}&\cdots&x_{1}^{T}Ax_{p}\\ \vdots&\ddots&\vdots\\ x_{p}^{T}Ax_{1}&\cdots&x_{p}^{T}Ax_{p}\end{pmatrix}$,
@@
Remark 4.3.
When AA is non-singular, Theorem 4.2 is a special case of Theorem 4.3. Notably, Theorem 4.2 remains applicable even when AA is singular. In contrast, Theorem 4.3 would require further modifications to accommodate the singular case, and we leave these extensions for future work.
@@ 4.4.2 $G$ is Unknown: Symmetry Discovery Framework
where $\hat{x}=Ux$, $U\in O(n)$ and $D$ is a diagonal matrix. In this formulation, $U$ captures the orthogonal transformation and $D$ scales the components of the input $x_{i}$, thereby enabling the discovery of the underlying symmetry group.

### v2

# Learning Equivariant Functions via Quadratic Forms

@@ 1 Introduction
Symmetry provides powerful inductive biases in deep learning, improving generalization and sample efficiency (Cohen, 2013; Cohen and Welling, 2014; Finzi et al., 2021; Dehmamy et al., 2021). When the underlying group is known, it can be encoded into architectures via invariant or equivariant operations, as in CNNs, which achieve translational equivariance through weight sharing.
Two main strategies exist for incorporating symmetries: (i) data augmentation with parameterized distributions (Benton et al., 2020), and (ii) architectures with intrinsic symmetry (Zaheer et al., 2017; Kicki et al., 2020; Zhou et al., 2020). In contrast, we develop a quadratic-form framework for learning $G$-equivariant functions, showing they canonically decompose into norm-invariant and scale-invariant terms.
Despite progress across images (Anselmi et al., 2019; Kondor and Trivedi, 2018), graphs (Han et al., 2022), and PDEs (Horie and Mitsume, 2023), challenges remain: most methods require prior group knowledge, rely on heavy computations (e.g., null-spaces, exponential maps), and scale poorly. Moreover, in many applications the group action is not directly observable, complicating symmetry exploitation.
We address these limitations by proposing a framework applicable with or without prior group knowledge. Focusing on orthogonal groups, we exploit their preservation of quadratic forms $x^{\top}Ax$. For $O(n)$ this corresponds to Euclidean norm preservation; for Lorentz groups it corresponds to the Minkowski norm, fundamental in physics. This geometric perspective enables efficient equivariance and symmetry discovery. Section 4 details the mathematical framework.
@@ 1.1 Contributions
  • We show that any function that is equivariant to a generalized orthogonal group (i.e., one that preserves a quadratic form $x^{T}Ax$) can be expressed as the product of two invariant functions: a scale-invariant component that acts on the input normalized by the quadratic-form norm, and a norm-invariant component that depends only on that norm. The product is interpreted in the context of the group’s action on the input space.
  • We extend this decomposition to the setting with diagonal group action on the input product space. In this case, the equivariant map splits into (i) an angular term computed from only the normalized first input, and (ii) a scale-invariant term that depends on the full Gram matrix (all pairwise quadratic-form inner products) of the inputs, thereby capturing their inter-dependencies.
  • Building on this decomposition, we introduce a symmetry discovery procedure: by learning the underlying quadratic form directly from data, the method uncovers the preserved symmetry without assuming it in advance.
  • We validate the approach across diverse tasks: synthetic polynomial regression, top-quark tagging (Lorentz-invariant classification), and prediction of moment-of-inertia matrices (rotation-equivariant regression), demonstrating accurate symmetry recovery and strong predictive performance.
@@ 2.2 Automatic Symmetry Discovery
Beyond architectures with fixed symmetries, several works focus on automatic discovery of symmetries directly from data. These approaches differ in how they parameterize transformations, enforce invariance or equivariance, and whether they yield a usable prediction model.
2.2.1 Distribution over generators
Augerino (Benton et al., 2020) learns a distribution over Lie algebra generators and samples transformations during training, applying an invariance or equivariance loss depending on the task. Similarly, LieGAN (Yang et al., 2023) learns such distributions but enforces consistency via an adversarial loss: a discriminator distinguishes whether transformed samples remain within the data distribution. A key difference is that Augerino yields a trained prediction model with soft equivariance, while LieGAN focuses on generator discovery and does not provide a predictor.
2.2.2 Post-hoc generator extraction
LieGG (Moskalev et al., 2022) extracts Lie algebra generators from trained models using linear-algebraic analysis. While useful for diagnosing symmetries, the approach is inherently post-hoc and does not construct an explicit equivariant model.
2.2.3 Vector-field approaches
Learning Infinitesimal Generators (LIG) (Ko et al., 2024) models continuous flows via Neural ODEs, extending symmetry discovery beyond affine groups. However, enforcing equivariance would require synchronized flows in the output space, computationally non-trivial and generally infeasible. Symmetry Discovery Beyond Affine (SDBA) (Shaw et al., 2024) also identifies vector fields that annihilate functions, but is typically limited to single-generator settings and mainly yields invariants rather than full equivariant predictors. Both methods therefore provide only indirect or implicit symmetry information.
2.2.4 Discrete-group inference
Bispectral Neural Networks (BNN) (Sanborn et al., 2022) address discrete groups by reconstructing Cayley tables from bispectrum features, thus inferring the group explicitly. While highly interpretable for finite groups, no analogous bispectral construction exists for continuous symmetries.
2.2.5 Our distinction
In contrast, our framework simultaneously discovers the underlying symmetry and provides an explicitly interpretable prediction model. By learning the quadratic form preserved by an orthogonal group, we obtain a canonical decomposition into norm- and scale-invariant components. Unlike distribution-based methods (Augerino, LieGAN), which enforce equivariance only implicitly, and unlike post-hoc or indirect approaches (LieGG, LIG, SDBA), our method integrates discovery and prediction in a unified and interpretable manner, combining the clarity of BNN with applicability to continuous groups.
@@ 3 Preliminaries
A group $(G,\cdot)$ is a set $G$ with a binary operation $\cdot:G\times G\to G$ satisfying: (i) Closure: $a\cdot b\in G$ for all $a,b\in G$; (ii) Associativity: $(a\cdot b)\cdot c=a\cdot(b\cdot c)$; (iii) Identity: $\exists e\in G$ s.t. $e\cdot a=a\cdot e=a$ for all $a\in G$; (iv) Inverse: $\forall a\in G,\ \exists a^{-1}\in G$ with $a\cdot a^{-1}=a^{-1}\cdot a=e$.
@@
A (left) group action of $G$ on a set $X$ is a map $G\times X\to X,\ (g,x)\mapsto g\cdot x$, such that (i) $e\cdot x=x$ for all $x\in X$; (ii) $(gh)\cdot x=g\cdot(h\cdot x)$ for all $g,h\in G,\ x\in X$.
@@
Given a group action of $G$ on $X$, the orbit of $x\in X$ is
@@
The general linear group $\mathrm{GL}(n,\mathbb{R})$ is the group of all invertible $n\times n$ real matrices under matrix multiplication.
@@
Let $G$ act on sets $X$ and $Y$. A function $f:X\to Y$ is equivariant if
$f(g\cdot x)=g\cdot f(x),\quad\forall g\in G,\ x\in X.$
Equivariance means $f$ commutes with the group action. An invariant function is the special case where $g\cdot f(x)=f(x)$ for all $g\in G$.
@@ 4 Proposed Method
4.1 Quadratic Forms
A quadratic form in $x\in\mathbb{R}^{n}$, with $A\in\mathbb{R}^{n\times n}$, is
$q(x;A):=x^{T}Ax.$ (2)
Every quadratic form is uniquely determined by a symmetric matrix, since
$q(x;A)=x^{T}Ax=\tfrac{1}{2}\,x^{T}(A+A^{T})x,\quad\forall x\in\mathbb{R}^{n},$
so without loss of generality we assume $A$ symmetric.
A symmetric matrix $A$ can be diagonalized by an orthogonal matrix $U$, i.e.
@@ 4.1 Quadratic Forms
where $D$ is diagonal with the eigenvalues of $A$. Substituting into (2) gives
$q(x;A)=x^{T}U^{T}DUx=(Ux)^{T}D(Ux)=y^{T}Dy,$ (4)
with $y=Ux$. Hence, learning $q(x;A)$ reduces to learning an orthogonal matrix $U$ and a diagonal matrix $D$.
4.2 Orthogonal Group: $O(p,q)$
Consider the set defined as,
$O(p,q):=\{\,U\in GL(n,\mathbb{R}):U^{T}AU=A\,\},$ (5)
which is a Lie subgroup of $GL(n,\mathbb{R})$ that preserves the quadratic form $q(x;A)=x^{T}Ax$, i.e. $q(Rx;A)=q(x;A)$ for all $R\in O(p,q)$. Examples include the standard orthogonal group preserving the Euclidean norm $x^{T}x$, and the Lorentz group preserving the Minkowski norm $x^{T}\eta x$ with $\eta=\mathrm{diag}(1,-1,-1,-1)$.
@@ 4.2 Orthogonal Group: $O(p,q)$
Remark 4.1.
We would like to emphasize that $\|z\|_{A}$ is a pseudonorm, as it can take on negative values.
@@ 4.3 Diagonal group action
Let $A$ be a non-singular $n\times n$ real matrix and define Let $G$ be orthogonal group as defined in eq. (5). Let $X\subset\mathbb{R}^{n}$ be a $G$-invariant set. For any positive integer $p$, consider the diagonal action of $G$ on the product space $X_{p}:=\underbrace{X\times X\times\cdots\times X}_{p\text{ times}},$ defined by, $\forall g\in G$
$g\cdot(x_{1},\dots,x_{p})=(g\cdot x_{1},\,\dots,\,g\cdot x_{p}).$
Then every $G$-equivariant function $f:X_{p}\to\mathbb{R}^{m}$ can be decomposed as
@@
$\operatorname{Gram}(x_{1},\ldots,x_{p})=\begin{bmatrix}x_{i}^{T}Ax_{j}\end{bmatrix}_{i,j\in[p]}$
@@
Remark $4.$
When $A$ is non-singular, Theorem 4.2 is a special case of Theorem 4.3. Notably, Theorem 4.2 remains applicable even when AA is singular. In contrast, Theorem 4.3 would require further modifications to accommodate the singular case, and we leave these extensions for future work.
@@ 4.4.2 $G$ is Unknown: Symmetry Discovery Framework
where $\hat{x}=Ux$, $U\in O(n)$ and $D$ is a diagonal matrix. In this formulation, $U$ captures the orthogonal transformation and $D$ scales the components of the input $x_{i}$, thereby enabling the discovery of the underlying symmetry group. The block diagram of our proposed method is depcited in Figure 1.